OpenStudy (anonymous):

If f(x) = 2x2 − 10, find f(5).

OpenStudy (anonymous):

Put x = 5: \[\huge \color{blue} {f(5) = 2(5)^2 - 10}\] Solve now..

OpenStudy (anonymous):

40

OpenStudy (anonymous):

Yes dear..

hero (hero):

Replace 5 everywhere you see x \(f(x)=2x^2-10\) \(f(5)=2(5)^2-10\)

hero (hero):

since x = 5

OpenStudy (anonymous):

10

OpenStudy (anonymous):

No dear you were right earlier...

hero (hero):

Dear?

OpenStudy (anonymous):

friend, buddy etc etc..

OpenStudy (anonymous):

So, what is the value of f(x) you got @ivandelgado6 ??

OpenStudy (anonymous):

its 10

OpenStudy (anonymous):

So how did you calculate 40 earlier??

OpenStudy (anonymous):

What is the square of 5??